OpenStudy (anonymous):

can someone decompose this into partial fractions? x-7/x(x^2+2)=A/x+Bx+C/x^2+2

OpenStudy (anonymous):

that looks right

OpenStudy (anonymous):

i need to solve for A, B, and C to get the answer

OpenStudy (anonymous):

hmmm. set x equal to zero and solve and see if a variable gives a number.

OpenStudy (anonymous):

Does the initial problem look like this? \[\frac{x-7}{x(x^2+2)}=\frac{A}{x}+Bx+\frac{C}{x^2+2}\]

OpenStudy (anonymous):

i want to say that the Bx+C should all be on top of the X^2+2

OpenStudy (anonymous):

Oh good, that makes more sense. \[\frac{x-7}{x(x^2+2)}=\frac{A}{x}+\frac{Bx+C}{x^2+2}\]Find a common denominator, namely x(x^2+2) \[\frac{x-7}{x(x^2+2)}=\frac{A(x^2+2)}{x(x^2+2)}+\frac{(Bx+C)x}{x(x^2+2)}\] Clear out the denominator. \[x-7=Ax^2+2A+Bx^2+Cx\] So here's what you have. You have a 0x^2 term, a 1x term, and a -7 term. \[x^2(A+B)=0x^2\]\[Cx=x\]\[2A=-7\] These equations yield \[A=-\frac{7}{2}, B=\frac{7}{2}, C=1\] Double-check my math, but I think I did that right...